# Degeneration of genus three Ceresa cycles and limit of archimedean height

### 02 Apr 2026 v1

@@ 4.2. Degeneration of height pairings
as a morphism of mixed Hodge structures. It is easy to see that $f(1)=1_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$ and $f(1^{\vee})=1^{\vee}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$, where we recall that $1_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$ is given by $\{\widetilde{Z}_{0}\}$, while $1^{\vee}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$ is given by the dual of $\{\widetilde{W}_{0}\}$. Since $f$ is a morphism of MHS, by Lemma 1.7 we have $f\circ\delta_{\mathcal{Q}}=\delta_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}\circ f$. Notice that since $B_{\widetilde{Z}_{0},\widetilde{W}_{0}}$ is a biextension MHS, we have $\delta_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}=\delta^{-1,-1}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$. On the other hand, $\mathcal{Q}$ is not a biextension mixed Hodge structure. But from the shape of its graded components we can conclude once again, using Lemma 1.7, and strictness of $f$, that $f\circ\delta^{-1,-1}_{\mathcal{Q}}=\delta^{-1,-1}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}\circ f$. Now we can invoke the argument of Proposition 3.4 in this situation to conclude
$H(N,F_{\infty},W)=\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}}).$
The case when $C_{0}=C_{1}\cup C_{2}$ is reducible follows the same set of ideas. In this scenario, we have $H^{1}_{\lim}(C_{t};\mathds{Q})\cong H^{1}(C_{1};\mathds{Q})\oplus H^{1}(C_{2};\mathds{Q})$, which is a pure Hodge structure and is automatically split. In this case, the biextension variation $B_{Z_{t},W_{t}}=H^{3}(J(C_{t})\setminus|Z_{t}|,|W_{t}|;\mathds{Q}(2))$ degenerates to a biextension mixed Hodge structure $B_{Z_{0},W_{0}}=H^{3}(X_{0}\smallsetminus|Z_{0}|,|W_{0}|;\mathds{Q}(2))$, where $X_{0}=JC_{1}\times JC_{2}$, $N=0$, and $M=W$. Hence, $\delta=\delta_{M}=\delta_{M,-2}$, and therefore
$\lim_{t\to 0}\operatorname{ht}(B_{Z_{t},W_{t}})=\operatorname{ht}(B_{Z_{0},W_{0}}).$
Notice that, in this case $Z_{0}$ and $W_{0}$ are homologous to zero, but may not be rationally equivalent to zero (Proposition 4.4). ∎

### 15 Apr 2026 v2

@@ 4.2. Degeneration of height pairings
as a morphism of mixed Hodge structures. It is easy to see that $f(1_{t})=1_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$ and $f(1^{\vee}_{t})=1^{\vee}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$, where we recall that $1_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$ is given by $\{\widetilde{Z}_{0}\}$, while $1^{\vee}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$ is given by the dual of $\{\widetilde{W}_{0}\}$. Since $f$ is a morphism of MHS, by Lemma 1.7 we have $f\circ\delta_{\mathcal{Q}}=\delta_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}\circ f$. Notice that since $B_{\widetilde{Z}_{0},\widetilde{W}_{0}}$ is a biextension MHS, we have $\delta_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}=\delta^{-1,-1}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$. On the other hand, $\mathcal{Q}$ is not a biextension mixed Hodge structure. But from the shape of its graded components we can conclude once again, using Lemma 1.7, and strictness of $f$, that $f\circ\delta^{-1,-1}_{\mathcal{Q}}=\delta^{-1,-1}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}\circ f$. Now we can invoke the argument of Proposition 3.4 in this situation to conclude
$\delta^{-1,-1}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}(1_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}})=\frac{1}{2\pi}H(N,F_{\infty},W)1^{\vee}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}.$
As an immediate consequence, we get $H(N,F_{\infty},W)=\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}})$.
The case when $C_{0}=C_{1}\cup C_{2}$ is reducible follows the same set of ideas. In this scenario, we have $H^{1}_{\lim}(C_{t};\mathds{Q})\cong H^{1}(C_{1};\mathds{Q})\oplus H^{1}(C_{2};\mathds{Q})$, which is a pure Hodge structure and is automatically split. Hence, unlike the previous scenario, a choice of a coordinate cannot be made at this stage. In this case, the biextension variation $B_{Z_{t},W_{t}}=H^{3}(J(C_{t})\setminus|Z_{t}|,|W_{t}|;\mathds{Q}(2))$ extends to a biextension mixed Hodge structure $B_{Z_{0},W_{0}}=H^{3}(X_{0}\smallsetminus|Z_{0}|,|W_{0}|;\mathds{Q}(2))$, where $X_{0}=JC_{1}\times JC_{2}$, and $M=W$. Notice that what we really have is the following: For any coordinate $t\in\Delta$, the limit of the biextension variation $B_{Z_{t},W_{t}}$ is isomorphic to $B_{Z_{0},W_{0}}$. However, this isomorphism is not fixed and depends on $t$. In this case, one can choose a coordinate $t$ in such a way that (as before) $1_{t}\mapsto\{Z_{0}\}$. In any case, we have that $H(N,F,W)$ is given by the Deligne $\delta$ corresponding to $B_{Z_{0},W_{0}}$. Notice that, in this case, $Z_{0}$ and $W_{0}$ are homologous to zero, but may not be rationally equivalent to zero (Proposition 4.4). ∎